OpenStudy (frostarmy):

Which shows -5x + 10y = 80 in slope-intercept form? Options: A. y = -2x + 8 B. y = 2x + 8 C. y = -1/2x + 8 D. y = 1/2x + 8

Vocaloid (vocaloid):

start by dividing both sides by 10

Vocaloid (vocaloid):

what is -5x/10?

OpenStudy (frostarmy):

-0.5

Vocaloid (vocaloid):

not quite

Vocaloid (vocaloid):

don't forget the x

OpenStudy (frostarmy):

oh -0.5x

Vocaloid (vocaloid):

good! how do we write that as a fraction?

OpenStudy (frostarmy):

1/2x

Vocaloid (vocaloid):

good what about 10y/10?

OpenStudy (frostarmy):

1y

Vocaloid (vocaloid):

good! we can also re-write 1y as y

Vocaloid (vocaloid):

80/10 = ?

OpenStudy (frostarmy):

8

Vocaloid (vocaloid):

good right now we have -(1/2)x + y = 8 we can add (1/2)x to both sides to get y = (1/2)x + 8 which answer choice matches?

OpenStudy (frostarmy):

D. y=(1/2)x + 8

Vocaloid (vocaloid):

good job!

OpenStudy (frostarmy):

ahhh thank you @Vocaloid